Update from Saturday 28th Feb and more recent photos!  Pics by Ian Robins and John Rees

D1661 – The fuel injectors were refitted into the 47 on Saturday 28/2.  
Whilst Neil, John Cooke and myself installed the overhauled units and associated pipework into the heads,
Cameron, Leroy and Dave L steam cleaned the rocker and fuel gallery covers.  Dave also replaced two
damaged rocker cover seals that should stem a couple of oil leaks.
Once complete, the engine was started and checks performed to ensure that there was no evidence of fuel
leakage into the oil, all was deemed ok.
The engine runs noticeably smoother with the new injectors but an oil change has to be performed before it
can re-enter service.


D6566 – I went to Cranmore on Sunday primarily to look at the marker light wiring in No.2 cab now they have
been made operational again by CTMS.   
On D6575 the wiring had been mostly removed with only a few inches left at the lamp ends which I spliced the
new cabling to.
I expected to find a similar situation on D6566 but when I removed the covers from the top of the lamp holders
the wiring was still intact!
The problem now was to find where the wires actually ended up as they disappeared inside the conduit so
were not physically possible to trace and are not included on the latest version of the wiring diagram.
According to the wiring diagram they were in fact isolated as long ago as 1973!
To cut a long story short, after some continuity testing I managed to find the original wiring to the switch
stowed up underneath the desk. These were both still tagged and idented and would have gone to the
original switch.
Judging by the cable length that would either have been located either in the main switch/instrument panel or
in the head code panel.
Back on the under desk terminal block, I found a tagged LP3 (Lighting Positive 3) cable taped and stowed
which rang to one of the switch wires so reconnected it in the appropriate position.
The other switch cable rang back to one pin on each of the lamp holders with the other pins going to LN
(Lighting Negative).
I had a look in No.1 cab and indications are that a similar situation exists although I couldn’t search for the
stowed cables as there was too much stuff stored on top of the desk.
So with new switches and some bulbs the marker lights should be back in business!

Other recent work was noted:

-Repairs to the air horn piping have been performed.

-Underneath the loco the new sump cover has been manufactured based on the original drawings.

-The cab steps are now being dealt with and differing amounts of repair work are required varying from
basically nothing to almost full rebuild!

-But aluminium tread plates will again be used in place of the troublesome wooded treads similar to D6575.

-The DSD pedal boxes are being painted and rebuilt ready to be re-fitted hence why I had to look at the
marker light wiring!

-I also checked that all the rear axle box covers are still fixed ok following recent problems found on another
Class 33. D6575 was similarly checked a couple of weeks ago and also found to be in order.